qually! Solve the following pair of linear equations; 8x + 6y = 12, y = - 4x/3 +2
Answer #1

8x + 64 = 12 Y = - 4x + 2 2y = - 4x + he 4x + 2 y = - 2 write in matrix form 101-13) argumented matrix (=[A:+] = 18 6:12] 143